OpenStudy (anonymous):

A solution is made by mixing equal masses of methanol, CH4O, and ethanol, C2H6O. Determine the mole fraction of each component to at least three significant figures.

OpenStudy (abmon98):

ethanol = 46.06844 g/mol Molar mass methanol = 32.04 Assume that you used 78.1044g of each: mol ethanol = 78.10844/46.06844 = 1.695486975 mol methanol = 78/32.04 = 2.437841448 Total mol = 1.695486975+2.437841448=4.133328423 mol fraction ethanol = 1.7/ 4.1 = 0.410198949 = 0.410 to 3 sig figs mol fraction methanol = 3.12110/4.133328423 = 0.589801051 = 0.590 to 3 sig figs